OpenStudy (anonymous):

3x^2-x-4, factor completely. <- HELP

OpenStudy (anonymous):

Please help.

OpenStudy (campbell_st):

ok... here is a method that works every time for a quadratic \[ax^2 + bx + c\] in your question a = 3 and c = -4 what is 3 x -4 = ? the find the factors that add to b, -1

OpenStudy (anonymous):

So I would have to multiply 3 time 4 which gets me 12. The factors would be 1, 2, 6, and 12. None that I see add to -1. :\ @campbell_st

OpenStudy (anonymous):

Can you tell me the factors of -4?

OpenStudy (campbell_st):

well 3 * -4 = -12 so find the factors that add to -1, the larger factor is negative...

OpenStudy (campbell_st):

then its simply a case of \[\frac{(ax + factor~1)(ax + factor~2)}{a}\]

OpenStudy (campbell_st):

and you missed 2 factors of 12... 3 and 4 so if you use -4 and 3 they add to -1 and multiply to -12 so you'll get \[\frac{(3x + 3)(3x -4)}{3}\] can you see a common factor in the 1st binomial that cancels with the denominator?

OpenStudy (campbell_st):

just remove the common factor for the answer

OpenStudy (anonymous):

Try to set up like this: (3x )(x ) Because of the -4, one factor has negative sign and other must have positive sign. So it can be like this (3x-#)(x+#) or (3x+#)(x-#)

OpenStudy (anonymous):